OpenStudy (anonymous):

How do i integrate this by parts x^11sin(x^6)dx?

OpenStudy (anonymous):

\[\int x^{11}Sin(x^6)dx\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

U sub u= x^6 du = 6x^5

OpenStudy (anonymous):

du* u= 6 x^11

OpenStudy (anonymous):

\[\frac{1}{6}\int u Sin(u) du \]

OpenStudy (anonymous):

do integration by parts a = u db= sin(u) da= 1 du b= -Cos(u)

OpenStudy (anonymous):

-u Cos(u)+ \[\int Cos(u) du\] \[\frac{1}{6}(-u Cos(u)+ Sin(u)+C)\]

OpenStudy (anonymous):

u= x^6 thus put that back in \[\frac{1}{6}(-x^6 Cos(x^6)+ Sin(x^6)+C)\]

OpenStudy (anonymous):

how did you get (1/6)?

OpenStudy (anonymous):

du* u= 6 x^11 we want x^11 so we divide by 6